OpenStudy (anonymous):

What is X? x/x-6 = 4/7

OpenStudy (anonymous):

I assume the left side is x/(x-6) Rearrange to get 7/4=(x-6)/x the right side can now be decomposed to x/x - 6/x which is equal to 1 - 6/x hence 7/4 = 1 - 6/x rearranging yields 6/x = 1 - 7/4 taking the reciprocal of both sides gives x/6 = 1/(1- 7/4) => x = 6/(1 - 7/4) = 6/(-3/4) = -(4*6)/3 = -8

OpenStudy (anonymous):

-8 is right!